# Torsion and Chern-Simons gravity in 4D space-times from a Geometrodynamical four-form

### v2

@@
Abstract
In Hermann Minkowski’s pioneering formulation of special relativity, the space-time geometry in any inertial frame is described by the line-element $ds^{2}=\eta_{\mu\nu}dx^{\mu}dx^{\nu}$. It is interesting to note that not only the Minkowski metric $\eta_{\mu\nu}$ is invariant under proper Lorentz transformations, the totally antisymmetric Levi-Civita tensor $e_{\mu\nu\alpha\beta}$ too is.
@@ Abstract
In Einstein’s general relativity (GR), $\eta_{\mu\nu}$ of the flat space-time gets generalized to a dynamical, space-time dependent metric tensor $g_{\mu\nu}$ that characterizes a curved space-time geometry. In the present study, it is put forward that the flat space-time Levi-Civita tensor gets elevated to a dynamical four-form field $\tilde{w}$ in curved space-time manifolds, i.e. $e_{\mu\nu\alpha\beta}\rightarrow w_{\mu\nu\alpha\beta}(x)=\phi(x)\ e_{\mu\nu\alpha\beta}$ so that $\tilde{w}={1\over{4!}}\ w_{\mu\nu\rho\sigma}\ \tilde{d}x^{\mu}\wedge\tilde{d}x^{\nu}\wedge\tilde{d}x^{\rho}\wedge\tilde{d}x^{\sigma}$. It is shown that this geometrodynamical four-form field, extends GR by leading naturally to a torsion in the theory as well as to a Chern-Simons gravity. Furthermore, it is demonstrated that the scalar-density $\phi(x)$ associated with the geometrodynamical four form $\tilde{w}$ may be used to construct a generalized exterior derivative that converts a p-form density to a (p+1)-form density of identical weight.
In order to subject the hypothesized four-form field $\tilde{w}$ to observational evidence, we first argue that the associated scalar-density $\phi(x)$ corresponds to an axion-like pseudo-scalar field in the Minkowski space-time, and that it can also masquerade as dark matter. Thereafter, we provide a simple semi-classical analysis in which a self-gravitating Bose-Einstein condensate of such ultra-light pseudo-scalars leads to the formation of a supermassive black hole. A brief analysis of propagation of weak gravitational waves in the presence of $\tilde{w}$ is also considered in this article.
I I. Introduction
@@
By envisaging a four dimensional manifold’s space-time geometry be described by a line-element, $ds^{2}=\eta_{\mu\nu}dx^{\mu}dx^{\nu}$, Hermann Minkowski had overturned the idea of treating space and time disjointly. Furthermore, the cause and effect relation, in the realm of classical physics, attained a significant elucidation in the Minkowskian framework - infinitesimally separated events (assuming a signature (+ - - -)) for which $ds^{2}<0$ ($ds^{2}>0$) are causally disconnected (causally connected). Of course, the underlying physics involved in it is that an effect essentially follows a cause due to transfer of energy from the latter to the former, and that the rate of flow of energy is limited by the speed $c$. What is usually not stressed is that, because of this causality criteria, the imaginary numbers entered physics in a fundamental manner even before the advent of quantum mechanics, i.e. the proper distance $ds$ between any two causally disconnected events is always an imaginary number.
@@
Although space and time coordinates get mixed up while transiting from one inertial frame to another, the Minkowski metric $\eta_{\mu\nu}$ itself remains invariant. In fact, this is utilized in the Minkowskian framework to define an arbitrary Lorentz transformation $x^{\mu}\rightarrow x^{{}^{\prime}\mu}=\Lambda^{\mu}_{\ \nu}x^{\nu}$ by the condition,
$\eta_{\mu\nu}=\Lambda^{\alpha}_{\ \mu}\Lambda^{\beta}_{\ \nu}\ \eta_{\alpha\beta}$ (1)
@@
$\epsilon_{\mu\nu\alpha\beta}\ \ \ \ \rightarrow\ \ \Lambda^{\sigma}_{\ \mu}\Lambda^{\tau}_{\ \nu}\Lambda^{\gamma}_{\ \alpha}\Lambda^{\kappa}_{\ \beta}\ \epsilon_{\sigma\tau\gamma\kappa}$ (2)
$=det(\Lambda)\ \epsilon_{\mu\nu\alpha\beta}\ \ ,$ (3)
@@
Given this special status that the Minkowski metric and the Levi-Civita tensor enjoy in the Minkowskian space-time, as well as the fact that $\eta_{\mu\nu}$ metamorphoses into a dynamical space-time metric $g_{\mu\nu}(x)$ in GR, raises a pertinent point of of elevating $\epsilon_{\mu\nu\alpha\beta}$ into a dynamical field. In the present article, we explore such a possibility by positing a geometrodynamical field, $w_{\mu\nu\rho\sigma}(x)$, totally antisymmetric in its indices, that is independent of the metric tensor and that influences space-time physics. Then, the interesting feature that $\eta_{\mu\nu}$ and $\epsilon_{\mu\nu\alpha\beta}$ both share in the flat space-time becomes relevant also in arbitrary curved space-times pdg-ref-20 .
We propose in this article that this geometrodynamical field represents a new physical degree of freedom that couples universally to all matter. In the following section, we discuss some of the differential geometric aspects of $w_{\mu\nu\rho\sigma}(x)$ that may be significant in the context of space-time physics.
II II. Geometrodynamical four-form field and Torsion
Now, in any 4D space-time manifold, because of its complete antisymmetry, $w_{\mu\nu\rho\sigma}(x)$ has only one algebraically independent component and, hence, renders itself to be expressed as,
$w_{\mu\nu\rho\sigma}=\phi(x)\epsilon_{\mu\nu\rho\sigma}\ \ ,$ (4)
@@
$\tilde{w}={1\over{4!}}\ w_{\mu\nu\rho\sigma}(x)\ \tilde{d}x^{\mu}\wedge\tilde{d}x^{\nu}\wedge\tilde{d}x^{\rho}\wedge\tilde{d}x^{\sigma}$ (5)
$={1\over{4!}}\ \phi(x)\ \epsilon_{\mu\nu\rho\sigma}\ \tilde{d}x^{\mu}\wedge\tilde{d}x^{\nu}\wedge\tilde{d}x^{\rho}\wedge\tilde{d}x^{\sigma}$ (6)
Eq.(6) reflects the equivalence between four-forms and 0-forms in the case of four dimensional manifolds.
Since we demand that this new dynamical field be independent of the metric tensor, instead of raising the indices of $w_{\mu\nu\rho\sigma}$ using the metric tensor, the totally antisymmetric components $w^{\mu\nu\rho\sigma}$ corresponding to the four-form $\tilde{w}$ are obtained from the associated p-vector w, so that one obtains the relation, $w^{\mu\nu\rho\sigma}\ w_{\mu\nu\rho\sigma}=-4!$ Schutz80 . This implies,
$w^{\mu\nu\rho\sigma}=\epsilon^{\mu\nu\rho\sigma}/\phi(x)\ \ .$ (7)
We assume throughout that $\phi=w_{0123}$ has no physical dimension. As the canonical volume-form in GR is given by,
$\tilde{V}={1\over{4!}}\sqrt{-g}\ \ \epsilon_{\mu\nu\rho\sigma}\ \tilde{d}x^{\mu}\wedge\tilde{d}x^{\nu}\wedge\tilde{d}x^{\rho}\wedge\tilde{d}x^{\sigma}\ ,$ (8)
our geometrodynamical four-form field may be expressed in terms of a scalar field $\chi(x)\equiv\frac{\phi(x)}{\sqrt{-g}}$ entailing,
$\tilde{w}=\chi(x)\tilde{V}$ (9)
A fundamental four-form field that varies with space and time may be utilized to extend GR by having a dynamical torsion field in the theory. Introducing an asymmetric affine connection (e.g. see Hehl ; Hehl1 ),
$\bar{\Gamma}^{\mu}_{\alpha\beta}\equiv\Gamma^{\mu}_{\alpha\beta}-a_{T}\ g^{\mu\nu}g^{\gamma\lambda}\frac{\partial\chi}{\partial x^{\lambda}}\ w_{\nu\gamma\alpha\beta}=\Gamma^{\mu}_{\alpha\beta}-a_{T}\ g^{\mu\nu}\chi^{\ ;\gamma}\ w_{\nu\gamma\alpha\beta}\ ,$ (10)
where $\Gamma^{\mu}_{\alpha\beta}$ is the standard Christoffel-Levi Civita connection while $a_{T}$ is a dimensionless constant that is a measure of the coupling between the metric tensor and the four-form, we define a metric compatible torsion field,
$S^{\ \ \ \mu}_{\alpha\beta}=\frac{1}{2}\bigg{(}\bar{\Gamma}^{\mu}_{\alpha\beta}-\bar{\Gamma}^{\mu}_{\beta\alpha}\bigg{)}=-a_{T}\ g^{\mu\nu}\chi^{\ ;\gamma}\ w_{\nu\gamma\alpha\beta}=-a_{T}\sqrt{-g}\ g^{\mu\nu}g^{\gamma\sigma}\ \chi\ \frac{\partial\chi}{\partial x^{\sigma}}\ \epsilon_{\nu\gamma\alpha\beta}$ (11)
@@
$D_{\nu}A^{\mu}=\partial_{\nu}A^{\mu}+\bar{\Gamma}^{\mu}_{\nu\beta}A^{\beta}=A^{\mu}_{\ ;\nu}-a_{T}\ g^{\mu\alpha}\chi^{\ ;\gamma}w_{\alpha\gamma\nu\beta}A^{\beta}\ \ ,$ (12)
$D_{\nu}B_{\mu}=\partial_{\nu}B_{\mu}-\bar{\Gamma}^{\beta}_{\mu\nu}B_{\beta}=B_{\mu;\ \nu}+a_{T}g^{\beta\sigma}\chi^{;\gamma}w_{\sigma\gamma\mu\nu}B_{\beta}\ \ ,$ (13)
@@
$D_{\nu}C_{\mu\tau}=\partial_{\nu}C_{\mu\tau}-\bar{\Gamma}^{\beta}_{\mu\nu}C_{\beta\tau}-\bar{\Gamma}^{\beta}_{\tau\nu}C_{\mu\beta}=C_{\mu\tau;\ \nu}+a_{T}g^{\beta\sigma}\chi^{;\gamma}w_{\sigma\gamma\mu\nu}C_{\beta\tau}+a_{T}g^{\beta\sigma}\chi^{;\gamma}w_{\sigma\gamma\tau\nu}C_{\mu\beta}\ \ ,$ (14)
with $C_{\mu\tau;\ \nu}$ being the covariant derivative of $C_{\mu\tau}$ in standard GR when there is no torsion (i.e. $a_{T}=0$ case). From eq.(14), the metric compatibility can be easily proved by showing that $D_{\alpha}g_{\mu\nu}=0$. Similarly, eq.(10) entails that,
@@
implying that the torsion in our case is autoparallel compatible. In our case, the contortion tensor $K^{\ \ \mu}_{\alpha\beta}$ and $S_{\alpha\beta\gamma}$ have the expressions,
$K^{\ \ \ \mu}_{\alpha\beta}\equiv-S^{\ \ \ \mu}_{\alpha\beta}+S^{\ \mu}_{\beta\ \alpha}-S^{\mu}_{\ \ \alpha\beta}=-S^{\ \ \ \mu}_{\alpha\beta}$ (15)
@@
$S_{\alpha\beta\gamma}=g_{\gamma\mu}S^{\ \ \ \mu}_{\alpha\beta}=a_{T}\ \chi^{\ ;\sigma}w_{\sigma\alpha\beta\gamma}=a_{T}\sqrt{-g}\ g^{\gamma\sigma}\chi\ \frac{\partial\chi}{\partial x^{\gamma}}\epsilon_{\sigma\alpha\beta\gamma}$ (16)
@@
III III. Action, Field Equations and Chern-Simons extensions
III.1 (i) Dynamical equations
@@
$\phi_{\ ;\beta}=\phi_{\ ,\beta}-\Gamma^{\alpha}_{\alpha\beta}\ \phi\ \ ,$
@@
$f_{\ ,\beta}\equiv\partial_{\beta}f\ \ .$
@@
$w_{\mu\nu\alpha\beta\ ;\lambda}=[(\ln\phi)_{,\lambda}-\Gamma^{\sigma}_{\sigma\lambda}]w_{\mu\nu\alpha\beta}\ \ ,$ (17)
@@
$w^{\mu\nu\alpha\beta}_{\ \ ;\lambda}=-[(\ln\phi)_{,\lambda}-\Gamma^{\sigma}_{\sigma\lambda}]w^{\mu\nu\alpha\beta}\ \ ,$ (18)
@@
$\ \ \ \ \ \ +{{a_{1}}\over{4!}}\int{\phi\ w^{\mu\nu\alpha\beta}_{\ \ ;\lambda}w_{\mu\nu\alpha\beta}\ ^{;\lambda}\ d^{4}x}+a_{2}\int{\phi(x)d^{4}x},$ (19)19( 19 )
where $m_{Pl}\equiv\sqrt{\hbar c/G}$ and $L$ are the Planck mass and the Lagrangian density of the matter fields, respectively, with $a_{1}$ and $a_{2}$ being real valued constants of the theory with dimensions $(\mbox{mass})^{2}$ and $(\mbox{mass})^{4}$, respectively. The part of the action in eq.(19) that pertains to the four-form is by no means unique. For instance, a term $\propto\int{R\ \phi d^{4}x}$ could be added to the above action, but we limit ourselves to gravitational minimal coupling, at present. Later, we shall discuss supplementing the above action with Chern-Simons terms induced by the geometrodynamical four-form.
By extremizing $\mathcal{A}$ with respect to $g_{\mu\nu}$ and $\phi$, respectively, we obtain the following equations of motion,
$R_{\mu\nu}-{1\over{2}}g_{\mu\nu}R={{8\pi}\over{m^{2}_{Pl}}}[T_{\mu\nu}+\Theta_{\mu\nu}],$ (20)20( 20 )
$\chi^{\ ;\alpha}_{\ ;\alpha}\equiv{1\over{\sqrt{-g}}}(\sqrt{-g}g^{\alpha\beta}\chi_{,\alpha})_{,\beta}={1\over{2}}\bigg{[}g^{\mu\nu}{{\chi_{,\mu}\chi_{,\nu}}\over{\chi}}+{a_{2}\over{a_{1}}}\chi\bigg{]},$ (21)21( 21 )
@@
$\Theta_{\mu\nu}=2a_{1}[{{\chi_{,\mu}\chi_{,\nu}}\over{\chi}}-g_{\mu\nu}\chi^{\ ;\alpha}_{\ ;\alpha}].$ (22)22( 22 )
When $\chi$ satisfies the equation of motion given by eq.(21), the above energy-momentum tensor gets simplified to,
$\Theta_{\mu\nu}=2a_{1}\bigg{[}{{\chi_{,\mu}\chi_{,\nu}}\over{\psi}}-{{g_{\mu\nu}}\over{2}}\bigg{(}g^{\alpha\beta}{{\chi_{,\alpha}\chi_{,\beta}}\over{\chi}}+{a_{2}\over{a_{1}}}\chi\bigg{)}\bigg{]}.$ (23)23( 23 )
III.2 (ii) Chern-Simons extensions
In this sub-section, we construct (3+1)-dimensional Chern-Simons (CS) terms. Our CS extensions of GR follows closely the seminal paper of Jackiw and Pi Jackiw , except for a crucial difference: they had introduced an external fixed four-vector $v_{\mu}$ instead of a dynamical $\phi_{;\mu}$ in their formulation. Because of this, their model had a manifest violation of Lorentz invariance. In our case, both general as well as Lorentz covariances are maintained throughout pdg-ref-20 .
First, we consider a Chern-Simons (CS) term that couples electromagnetic field to the four-form $\tilde{w}$,

### v3

# Patrick Das Gupta

@@
Abstract In Hermann Minkowski’s pioneering mathematical formulation of special relativity, the space-time geometry in any inertial frame is described by the line-element $ds^{2}=\eta_{\mu\nu}dx^{\mu}dx^{\nu}$. It is interesting to note that not only the Minkowski metric $\eta_{\mu\nu}$ is invariant under proper Lorentz transformations, the totally antisymmetric Levi-Civita tensor $e_{\mu\nu\alpha\beta}$ too is. In Einstein’s general relativity (GR), $\eta_{\mu\nu}$ of the flat space-time gets generalized to a dynamical, space-time dependent metric tensor $g_{\mu\nu}$ that characterizes a curved space-time geometry. In the present study, it is put forward that the flat space-time Levi-Civita tensor gets elevated to a dynamical four-form field $\tilde{w}$ in curved space-time manifolds, i.e. $e_{\mu\nu\alpha\beta}\rightarrow w_{\mu\nu\alpha\beta}(x)=\phi(x)\ e_{\mu\nu\alpha\beta}$, so that $\tilde{w}={1\over{4!}}\ w_{\mu\nu\rho\sigma}\ \tilde{d}x^{\mu}\wedge\tilde{d}x^{\nu}\wedge\tilde{d}x^{\rho}\wedge\tilde{d}x^{\sigma}$. It is shown that this geometrodynamical four-form field extends GR by leading naturally to a torsion in the theory as well as to a Chern-Simons gravity. Furthermore, it is demonstrated that the scalar-density $\phi(x)$ associated with the geometrodynamical four-form $\tilde{w}$ may be used to construct a generalized exterior derivative that converts a p-form density to a (p+1)-form density of identical weight. In order to relate the hypothesized four-form field $\tilde{w}$ to observational evidence, we first argue that the associated scalar-density $\phi(x)$ corresponds to an axion-like pseudo-scalar field in the Minkowski space-time, and that it can also masquerade as dark matter. Thereafter, we provide a simple semi-classical analysis in which a self-gravitating Bose-Einstein condensate of such ultra-light pseudo-scalars leads to the formation of a supermassive black hole. A brief analysis of propagation of weak gravitational waves in the presence of $\tilde{w}$ is also considered in this article.
Keywords: Torsion, Chern-Simons gravity, Dynamical four-form, Ultra-light Dark Matter, Supermassive Black holes, Bose-Einstein Condensation
1 Introduction
@@
By envisaging a four dimensional manifold’s space-time geometry be described by a line-element, $ds^{2}=\eta_{\mu\nu}dx^{\mu}dx^{\nu}$, Hermann Minkowski had overturned the idea of treating space and time disjointly. Furthermore, the cause and effect relation, in the realm of classical physics, attained a significant elucidation in the Minkowskian framework: infinitesimally separated events (assuming a signature (+ - - -)) for which $ds^{2}<0$ ($ds^{2}\geq 0$) are causally disconnected (causally connected). Of course, the underlying physics involved in it is that an effect essentially follows a cause due to transfer of energy from the latter to the former, and that the rate of flow of energy is limited by the speed $c$. What is usually not stressed is that, because of this causality criteria, the imaginary numbers entered physics in a fundamental manner even before the advent of quantum mechanics, i.e. the proper distance $ds$ between any two causally disconnected events is always an imaginary number.
@@
Although space and time coordinates get mixed up while transiting from one inertial frame to another, the Minkowski metric $\eta_{\mu\nu}$ itself remains invariant. This is utilized in the Minkowskian framework to define an arbitrary Lorentz transformation $x^{\mu}\rightarrow x^{{}^{\prime}\mu}=\Lambda^{\mu}_{\ \nu}x^{\nu}$ by the condition,
$\eta_{\mu\nu}=\Lambda^{\alpha}_{\ \mu}\Lambda^{\beta}_{\ \nu}\ \eta_{\alpha\beta}$ (1.1)
@@
$\epsilon_{\mu\nu\alpha\beta}\ \ \ \ \rightarrow\ \ \Lambda^{\sigma}_{\ \mu}\Lambda^{\tau}_{\ \nu}\Lambda^{\gamma}_{\ \alpha}\Lambda^{\kappa}_{\ \beta}\ \epsilon_{\sigma\tau\gamma\kappa}$ (1.2)
$=det(\Lambda)\ \epsilon_{\mu\nu\alpha\beta}\ \ ,$ (1.3)
@@
Given this special status that the Minkowski metric and the Levi-Civita tensor enjoy in the Minkowskian space-time, as well as the fact that $\eta_{\mu\nu}$ metamorphoses into a dynamical space-time metric $g_{\mu\nu}(x)$ in GR, raises a pertinent point of elevating $\epsilon_{\mu\nu\alpha\beta}$ to a dynamical field. In the present article, we explore such a possibility by positing a geometrodynamical field, $w_{\mu\nu\rho\sigma}(x)$, totally antisymmetric in its indices, that is independent of the metric tensor and that influences space-time physics. Then, the interesting feature that $\eta_{\mu\nu}$ and $\epsilon_{\mu\nu\alpha\beta}$ both share in the flat space-time becomes relevant also in arbitrary curved space-times [1].
We propose that this geometrodynamical field represents a new physical degree of freedom that couples universally to all matter. In the following section, we discuss some of the differential geometric aspects of $w_{\mu\nu\rho\sigma}(x)$ that may be significant in the context of space-time physics.
2 Geometrodynamical four-form field and Torsion
Now, in any 4D space-time manifold, because of its complete antisymmetry, the four-form field $w_{\mu\nu\rho\sigma}(x)$ has only one algebraically independent component and hence, renders itself to be expressed as,
$w_{\mu\nu\rho\sigma}(x)=\phi(x)\ \epsilon_{\mu\nu\rho\sigma}\ \ ,$ (2.1)
@@ 2 Geometrodynamical four-form field and Torsion
$\tilde{w}={1\over{4!}}\ w_{\mu\nu\rho\sigma}(x)\ \tilde{d}x^{\mu}\wedge\tilde{d}x^{\nu}\wedge\tilde{d}x^{\rho}\wedge\tilde{d}x^{\sigma}$ (2.2)
$={1\over{4!}}\ \phi(x)\ \epsilon_{\mu\nu\rho\sigma}\ \tilde{d}x^{\mu}\wedge\tilde{d}x^{\nu}\wedge\tilde{d}x^{\rho}\wedge\tilde{d}x^{\sigma}$ (2.3)
Eq.(2.3) reflects the equivalence between four-forms and 0-forms in the case of four dimensional manifolds.
Since we demand that this new dynamical field be independent of the metric tensor, instead of raising the indices of $w_{\mu\nu\rho\sigma}$ using the metric tensor, the totally antisymmetric components $w^{\mu\nu\rho\sigma}(x)$ corresponding to the four-form field $\tilde{w}$ are obtained from the associated p-vector w, so that one obtains the relation, $w^{\mu\nu\rho\sigma}\ w_{\mu\nu\rho\sigma}=-4!$ [2]. This implies,
$w^{\mu\nu\rho\sigma}(x)=\epsilon^{\mu\nu\rho\sigma}/\phi(x)\ \ .$ (2.4)
We assume throughout that $\phi(x)=w_{0123}(x)$ has no physical dimension. As the canonical volume-form in GR is given by,
$\tilde{V}={1\over{4!}}\sqrt{-g}\ \ \epsilon_{\mu\nu\rho\sigma}\ \tilde{d}x^{\mu}\wedge\tilde{d}x^{\nu}\wedge\tilde{d}x^{\rho}\wedge\tilde{d}x^{\sigma}\ ,$ (2.5)
the geometrodynamical four-form field of eq.(2.3) may be expressed in terms of a scalar field $\chi(x)$,
$\chi(x)\equiv\frac{\phi(x}{\sqrt{-g}}\ \ \Rightarrow\ \ \tilde{w}=\chi(x)\tilde{V}$ (2.6)
A fundamental four-form field that varies with space and time may be utilized to extend GR by having a dynamical torsion field in the theory. Introducing an asymmetric affine connection (e.g. see [3, 4]),
$\bar{\Gamma}^{\mu}_{\alpha\beta}\equiv\Gamma^{\mu}_{\alpha\beta}-a_{T}\ g^{\mu\nu}g^{\gamma\lambda}\frac{\partial\chi}{\partial x^{\lambda}}\ w_{\nu\gamma\alpha\beta}=\Gamma^{\mu}_{\alpha\beta}-a_{T}\ g^{\mu\nu}\chi^{;\gamma}\ w_{\nu\gamma\alpha\beta}\ ,$ (2.7)
where $\Gamma^{\mu}_{\alpha\beta}$ is the standard Christoffel-Levi Civita connection while $a_{T}$ is a dimensionless constant that is a measure of the coupling between the metric tensor and the four-form, we define a torsion field,
$S^{\ \ \ \mu}_{\alpha\beta}(x)=\frac{1}{2}\bigg{(}\bar{\Gamma}^{\mu}_{\alpha\beta}-\bar{\Gamma}^{\mu}_{\beta\alpha}\bigg{)}=-a_{T}\ g^{\mu\nu}\chi^{;\gamma}\ w_{\nu\gamma\alpha\beta}=-a_{T}\sqrt{-g}\ g^{\mu\nu}g^{\gamma\sigma}\ \chi\ \frac{\partial\chi}{\partial x^{\sigma}}\ \epsilon_{\nu\gamma\alpha\beta}$ (2.8)
@@
$D_{\nu}A^{\mu}=\partial_{\nu}A^{\mu}+\bar{\Gamma}^{\mu}_{\nu\beta}A^{\beta}=A^{\mu}_{\ ;\nu}-a_{T}\ g^{\mu\alpha}\chi^{;\gamma}w_{\alpha\gamma\nu\beta}A^{\beta}\ \ ,$ (2.9)
$D_{\nu}B_{\mu}=\partial_{\nu}B_{\mu}-\bar{\Gamma}^{\beta}_{\mu\nu}B_{\beta}=B_{\mu;\ \nu}+a_{T}g^{\beta\sigma}\chi^{;\gamma}w_{\sigma\gamma\mu\nu}B_{\beta}\ \ ,$ (2.10)
@@
$D_{\nu}C_{\mu\tau}=\partial_{\nu}C_{\mu\tau}-\bar{\Gamma}^{\beta}_{\mu\nu}C_{\beta\tau}-\bar{\Gamma}^{\beta}_{\tau\nu}C_{\mu\beta}=C_{\mu\tau;\ \nu}+a_{T}g^{\beta\sigma}\chi^{;\gamma}[w_{\sigma\gamma\mu\nu}C_{\beta\tau}+w_{\sigma\gamma\tau\nu}C_{\mu\beta}]\ \ ,$ (2.11)
with $C_{\mu\tau;\ \nu}$ being the covariant derivative of $C_{\mu\tau}$ in standard GR when there is no torsion (i.e. $a_{T}=0$ case). Using eq.(2.11), the metric compatibility of $S^{\ \ \ \mu}_{\alpha\beta}$ can easily be proved by showing that $D_{\alpha}g_{\mu\nu}=0$. Similarly, for any vector field $A^{\alpha}(x)$, eq.(2.7) entails that,
@@
implying that the torsion field given by eq.(2.8) is autoparallel compatible too. In our case, the contortion tensor $K^{\ \ \mu}_{\alpha\beta}$ and $S_{\alpha\beta\gamma}$ have the expressions,
$K^{\ \ \ \mu}_{\alpha\beta}\equiv-S^{\ \ \ \mu}_{\alpha\beta}+S^{\ \mu}_{\beta\ \alpha}-S^{\mu}_{\ \ \alpha\beta}=-S^{\ \ \ \mu}_{\alpha\beta}$ (2.12)
@@
$S_{\alpha\beta\gamma}=g_{\gamma\mu}S^{\ \ \ \mu}_{\alpha\beta}=a_{T}\ \chi^{;\sigma}w_{\sigma\alpha\beta\gamma}=a_{T}\sqrt{-g}\ g^{\gamma\sigma}\chi\ \frac{\partial\chi}{\partial x^{\gamma}}\epsilon_{\sigma\alpha\beta\gamma}$ (2.13)
@@
3 Action, Field Equations and Chern-Simons extensions
3.1 Dynamical equations
@@
$\phi_{\ ;\beta}=\phi_{\ ,\beta}-\Gamma^{\alpha}_{\alpha\beta}\ \phi\ \ ,$ (3.1)
@@
$f_{\ ,\beta}\equiv\partial_{\beta}f\ \ .$ (3.2)
@@
$w_{\mu\nu\alpha\beta\ ;\lambda}=[(\ln\phi)_{,\lambda}-\Gamma^{\sigma}_{\sigma\lambda}]w_{\mu\nu\alpha\beta}\ \ ,$ (3.3)
@@
$w^{\mu\nu\alpha\beta}_{\ \ ;\lambda}=-[(\ln\phi)_{,\lambda}-\Gamma^{\sigma}_{\sigma\lambda}]w^{\mu\nu\alpha\beta}\ \ ,$ (3.4)
@@
$\ \ \ \ \ \ +\ {{a_{1}}\over{4!}}\int{\phi\ w^{\mu\nu\alpha\beta}_{\ \ ;\lambda}w_{\mu\nu\alpha\beta}\ ^{;\lambda}\ d^{4}x}+a_{2}\int{\phi(x)d^{4}x}\ ,$ (3.5)
where $m_{Pl}\equiv\sqrt{\hbar c/G}$ and $L$ are the Planck mass and the Lagrangian density of the matter fields, respectively, with $a_{1}$ and $a_{2}$ being real valued constants of the theory having physical dimensions $(\mbox{mass})^{2}$ and $(\mbox{mass})^{4}$, respectively. The part of the action in eq.(3.5) that pertains to the four-form is by no means unique. For instance, a term $\propto\int{R\ \phi\ d^{4}x}$ could be added to the above action, but we limit ourselves to gravitational minimal coupling, at present. Later, we shall discuss supplementing the above action with Chern-Simons terms induced by the geometrodynamical four-form.
By extremising $\mathcal{A}$ with respect to $g_{\mu\nu}$ and $\phi$, respectively, we obtain the following equations of motion,
$R_{\mu\nu}-{1\over{2}}g_{\mu\nu}R={{8\pi}\over{m^{2}_{Pl}}}[T_{\mu\nu}+\Theta_{\mu\nu}]\ ,$ (3.6)
$\chi^{\ ;\alpha}_{\ ;\alpha}\equiv{1\over{\sqrt{-g}}}(\sqrt{-g}g^{\alpha\beta}\chi_{,\alpha})_{,\beta}={1\over{2}}\bigg{[}g^{\mu\nu}{{\chi_{,\mu}\chi_{,\nu}}\over{\chi}}+{a_{2}\over{a_{1}}}\chi\bigg{]}\ ,$ (3.7)
@@
$\Theta_{\mu\nu}=2a_{1}\bigg{[}{{\chi_{,\mu}\chi_{,\nu}}\over{\chi}}-g_{\mu\nu}\chi^{\ ;\alpha}_{\ ;\alpha}\bigg{]}\ .$ (3.8)
When $\chi$ satisfies the equation of motion given by eq.(3.7), the above energy-momentum tensor gets simplified to,
$\Theta_{\mu\nu}=2a_{1}\bigg{[}{{\chi_{,\mu}\chi_{,\nu}}\over{\psi}}-{{g_{\mu\nu}}\over{2}}\bigg{(}g^{\alpha\beta}{{\chi_{,\alpha}\chi_{,\beta}}\over{\chi}}+{a_{2}\over{a_{1}}}\chi\bigg{)}\bigg{]}\ .$ (3.9)
3.2 Chern-Simons extensions
In this sub-section, we construct Chern-Simons (CS) terms in (3+1)-dimensional space-time. Our CS extensions of GR follows closely the seminal paper of Jackiw and Pi [6], except for a crucial difference: they had introduced an external fixed four-vector $v_{\mu}$ instead of a dynamical $\phi_{;\mu}$ in their formulation. Because of this, their model had a manifest violation of Lorentz invariance. In our case, both general as well as Lorentz covariances are maintained throughout [1].
First, we consider a Chern-Simons (CS) term that couples the electromagnetic field to the four-form $\tilde{w}$,